# Distributed network of smartphone sensors: a new tool for scientific field measurements

### v2

@@ III.2 Wave buoys
The smartphone fleet can also be used as a network of autonomous IMU systems in outdoor conditions. We perform a proof of concept during a field campaign on wave sea ice interaction in Rimouski, Québec, Canada, during February 2024, where the smartphones were deployed as local wave buoys to record the motion of ice induced by gravity waves. Smartphones can aptly be used to record the motion of the ice induced by gravity waves, as an array of smartphones can provide spatial and temporal information on wave propagation and attenuation. Custom cases were 3D-printed to shield phones from the environmental condition, ensuring their temperature remains within working range.
@@
On the 23th of February, we noticed wave activity sufficient to break the sea ice cover (thickness of 16 cm) in the Mercier Cove, in the estuary of the Saint Lawrence river. We placed ten phones close to the forming ice edge, on a straight line, aligned with the direction of wave propagation, regularly spaced 10 meters apart. We recorded GPS coordinates, as well as acceleration, gyroscope and magnetometer signals. The recordings overlapped in time for about thirty minutes, and we only present the exploitation of vertical accelerations in this section, as a proxy for wave height. We show a top-view of the cropped situation, photographed by a drone, in Fig. 7(a).
@@
Two of these phones were doubled with wave buoys developed by P. Sutherland from Ifremer Sutherland and Dumont (2018), themselves equipped with geolocation and accelerometers, similar to the model described by Guimarães et al. Veras Guimarães et al. (2018). After the time-synchronization of the phones signals, we set their time origin by finding the lag maximizing the cross-correlation between the acceleration signals of the phone (#T19) and the buoy (#B1) situated at the same location. We show the excellent correlation between the signals in Fig. 7(b).
In Fig. 7(c), we show the time evolution of the vertical acceleration as recorded by phones at increasing distance from the ice edge. Qualitatively, we confirm the synchronization of the signals by observing individual wave packets moving in space. This recording reveals the attenuation of the wave amplitude as the wave propagates under partially fragmented ice.
IV Conclusion
@@
We show that a fleet of smartphone can be controlled from a single computer, to perform spatial measurements on mechanical vibrations. We first evaluate the sensors sensitivity and accuracy of a single smartphone and precisely determined the sensor locations. Based on existing protocols for single smartphones, we introduced a new architecture based on two softwares, Gobannos and Phonefleet. Gobannos is installed on each smartphone, and allow to remote control the entire acquisition process using url requests. Phonefleet is deployed on the controller, to send the instructions to the fleet, and gather the results. The combination of the two, with an appropriate local Wi-Fi network opens new avenue for smartphone-based multi-physics instrumentation. We eventually demonstrate the capabilities, by conducting experiments with the smartphone fleet in two model situations. The first experiment focuses on the large-scale physical measurements of turbulent fluctuations using smartphones as pendulums. The second experiment focuses on wave-ice floe interactions using smartphones as wave buoys. Altogether, this work paves the way for low cost and eco-friendly solutions for large-scale scientific studies.
Acknowledgements.
@@
We thank B. Auvity, S. Kuchly, B. Lafoux and G. Vandenhove for scientific discussions & technical developments, as well as PMMH members C. Aracheloff, R. Baillou, A. Bouvier, S. Gomé, N. He, S. Kumar Saroj, A. Mongruel, B. Musci, P. Petitjeans, W. Reino, A. Rivière, J. Tampier, V. Thievenaz, L. Tuckermann and D. Vandembroucq for providing their SIM card to unlock 3 smartphones of the fleet. We thank C. Robert and the technical staff of IAT Saint-Cyr-l’École for their help with the wind tunnel experiment. We also thank T-CAPES Initiative for scientific support. This work has benefited from the financial support of Mairie de Paris through Emergence(s) grant 2021-DAE-100 245973, of the Agence Nationale de la Recherche through grants MSIM ANR-23-CE01-0020-02 & LASCATURB ANR-23-CE30-0043-03, and of the PSL University through a Junior Fellowship 2022-305.

### v5

@@ 3.2 Wave buoys
The smartphone fleet can also be used as a network of autonomous IMU systems in outdoor conditions. We perform a proof of concept during a field campaign on wave-sea ice interaction in Rimouski, Québec, Canada, in February 2024, where the smartphones were deployed as local wave buoys to record the motion of ice induced by gravity waves [kuchly2025integrated]. Smartphones can aptly be used to record the motion of the ice induced by gravity waves, as an array of smartphones can provide spatial and temporal information on wave propagation and attenuation. Custom cases were 3D-printed to shield phones from the environmental condition, ensuring their temperature remains within working range.
@@
On 23 February, we noticed wave activity (thickness of 16 cm) in Mercier Cove, in the estuary of the Saint Lawrence river. We placed ten phones close to the forming ice edge, on a straight line, aligned with the direction of wave propagation, regularly spaced 10 meters apart. We recorded GPS coordinates, as well as acceleration, gyroscope and magnetometer signals. The recordings overlapped in time for about thirty minutes, and we only present the exploitation of vertical accelerations in this section, as a proxy for wave height. We show a top-view of the cropped situation, photographed by a drone, in Fig. 7(a).
@@
Two of these phones were doubled with wave buoys developed by P. Sutherland from Ifremer [Sutherland_2018], themselves equipped with geolocation and accelerometers, similar to the model described by Guimarães et al. [Guimaraes_2018]. After the time-synchronization of the phones signals, we set their time origin by finding the lag maximizing the cross-correlation between the acceleration signals of the phone (#T19) and the buoy (#B1) situated at the same location. We show an excellent agreement between the wave buoy signal and the smartphone accelerometer sensor (Fig. 7(b)). In Fig. 7(c), we show the time evolution of the vertical acceleration as recorded by 10 phones at increasing distance from the ice edge. Qualitatively, we confirm the synchronization of the signals by observing individual wave packets propagating in space. These recordings reveal the attenuation of the wave amplitude as the wave propagates under partially fragmented ice.
Figure 8: Measured in-ice wave propagation by phone accelerometers placed at increasing distances from the ice edge, in the direction of wave propagation. (a) Temporal spectra $\hat{a}_{z}$ of the vertical acceleration. The color encodes the distance to the ice edge. The peak at $f_{0}=0.3\pm 0.02$Hz corresponds to the most energetical waves. (b) Acceleration amplitude associated with wave-induced vertical motion inferred from the acceleration spectra, as a function of the distance from the ice edge (points), and exponential fit (dashed line) with an attenuation length $\ell_{c}=37.3$ m. (c) Zoom in view of the spectra normalised by the spectrum of the first phone $\hat{a}/\hat{a}_{0}$, showing the fast decay of the wave energy at higher frequency. (d) Measure of the attenuation coefficient $\alpha(f)$ per frequency, in the range $f\in[0.2,0.45]$Hz, extracted from an exponential fit of $\hat{a}/\hat{a}_{0}$ as a function of $L$. We found approximately $\alpha=\alpha_{0}f^{-2.2}$ with $\alpha_{0}=0.31$/m from a fit by a power law (red dashed line). Insert : Normalised acceleration amplitude to check the validity of the exponential decay model in the entire frequency range.
Fig. 8 shows the temporal Fourier spectrum of $a_{z}$ for the 10 smartphones, where the color corresponds to the distance to the ice edge. All the spectra present a maximum at $f_{0}=0.3\pm 0.02$Hz, and significant wave energy in the principal peak between 0.2 to 0.45 Hz, corresponding to the coastal swell propagating under the ice. Note that higher frequencies component of the spectrum are a sum of non linear contribution of the main peak, and energy at higher frequency. From these spectra, we compute a proxy of the wave amplitude $a_{w}$ as the standard deviation of $a_{z}$, filtered in the frequency range [0.1,5] Hz. The wave amplitude as a function of the ice edge is shown in Fig. 8b and is well fitted by an exponential decay, $a_{w}=a_{0}\textrm{exp}(-x/\ell_{c})$, where $x$ is the distance to the ice edge, and $\ell_{c}=37.3$m is a characteristic decay length. The wave amplitude $A$ can be inferred from the acceleration signal as $A~\sim 2(\pi f_{0})^{2}$, corresponding to $A$ ranging from 4cm at the ice edge, to 3mm at 90m from the ice edge. It is remarkable to measure the wave amplitude and the decay for such small wave amplitudes. We conducted additional analysis to extract a damping coefficient as a function of the wave frequency. Figure 8c) shows a zoom in view of the spectrum, normalised by the wave spectrum of the first phone located at the ice edge. From these spectrum, we fit an exponential decay $\hat{a}$ = $\hat{a}_{0}exp(-\alpha(f)L)$, and we extract the damping coefficient $\alpha$, shown in figure 8d as a function of the wave frequency. We checked that the exponential fit was a valid model, by representing $\alpha(f)\textrm{log}(\hat{a}_{z}/\hat{a}_{ref})$ as a function of the distance $L$ to the ice edge. The result is shown in the insert of figure 8d, where the color encodes the wave frequency. We found a good agreement with an exponential decay, all the curves lying in the vicinity of a linear segment of slope -1 (red dashed line). Unfortunately, the ice parameters could not be inferred from these measurements, as the range of excited frequency is too small to be sensitive to the ice properties (thickness, Young’s modulus, Poisson ratio).
These measurements show that smartphones can be deployed and used reliably to access wave data in outdoor conditions, and characterized typical properties of ocean waves, such as wave spectrum, or spatial decay. Although we did not manage to extract ice mechanical properties from these preliminary experiments, we believe that further works focusing on a higher frequency range combined with the use active sources, could provide an estimate of the Young’s modulus or the ice thickness from wave propagation.
4 Discussion
We have presented a methodology that transforms homogeneous smartphones into a calibrated, time-synchronised, and remotely coordinated large-scale sensor fleet for physical field measurements. By achieving microsecond-level synchronization, enabling centralised control and parallel data acquisition across tens of devices, this approach supports coherent multi-point measurements of rapidly evolving physical phenomena such as mechanical waves by wind turbulence or wave–ice interactions. Compared with established wireless sensor networks (WSNs) and IoT-based frameworks, our system overcomes several long-standing limitations: typical WSN synchronization protocols are constrained by heterogeneous devices and hardware clock-drift accumulation, which generally limit precision to the millisecond or tens-of-microseconds range. In contrast, our smartphone fleet benefits from homogeneous, well-characterised sensors, stable clocks, and a dedicated synchronization protocol that consistently achieves µs-level alignment, enabling coherent spatiotemporal reconstruction of dynamic fields. This positions our framework between ad-hoc smartphone sensing and specialised professional distributed instrumentation, offering the precision and cost-effectiveness required for dense field deployments. Overall, our results demonstrate that widely available smartphones can form a reliable scientific instrument for distributed measurements, opening new possibilities for large-scale, accessible field experiments in environmental, mechanical, and geophysical systems.
5 Conclusion
@@
We show that a fleet of smartphone can be controlled from a single computer, to perform spatial measurements of mechanical vibrations. We first evaluate the sensors sensitivity and accuracy for a single smartphone and precisely determine the sensor locations. Based on existing protocols for single smartphones, we introduced a new architecture based on two softwares, Gobannos and Phonefleet. Gobannos is installed on each smartphone, and allows to remote-control the entire acquisition process using url requests. Phonefleet is deployed on the controller, to send the instructions to the fleet, and gather the results. The combination of the two, with an appropriate local Wi-Fi network opens new avenue for smartphone-based multi-physics instrumentation. We eventually demonstrate the capabilities, by conducting experiments with the smartphone fleet in two model situations. The first experiment focuses on the large-scale physical measurements of turbulent fluctuations using smartphones as pendulums. The second experiment focuses on wave-ice floe interactions using smartphones as local wave buoys. Altogether, this work paves the way for low cost and eco-friendly solutions for large-scale scientific studies.
CRediT authorship contribution statement
Jishen Zhang: Conceptualization, Methodology, Investigation, Formal analysis, Writing - Original Draft. Nicolas Mokus: Investigation, Writing - Review & Editing. Jules Casoli: Software, Writing - Review & Editing. Antonin Eddi: Conceptualization, Methodology, Software, Investigation, Formal analysis, Writing - Original Draft. Stéphane Perrard: Conceptualization, Methodology, Software, Investigation, Formal analysis, Writing - Original Draft.
@@
acknowledgments We thank B. Auvity, S. Kuchly, B. Lafoux and G. Vandenhove for scientific discussions & technical developments, as well as PMMH members C. Aracheloff, R. Baillou, A. Bouvier, S. Gomé, N. He, S. Kumar Saroj, A. Mongruel, B. Musci, P. Petitjeans, W. Reino, A. Rivière, J. Tampier, V. Thievenaz, L. Tuckermann and D. Vandembroucq for providing their SIM card to unlock each 3 smartphones of the fleet. We thank C. Robert and the technical staff of IAT Saint-Cyr-l’École for their help with the wind tunnel experiment. We also thank T-CAPES Initiative for scientific and financial support. This work has benefited from the financial support of Mairie de Paris through Emergence(s) grant 2021-DAE-100 245973, of the Agence Nationale de la Recherche through grants MSIM ANR-23-CE01-0020-02 & LASCATURB ANR-23-CE30-0043-03, and of the PSL University through a Junior Fellowship 2022-305.
Appendix A Sensor calibration and accuracy tests
A.1 Gyroscope
Figure 9: Calibration of accelerometer and gyroscope. The measurement depends on the axis of rotation. When the smartphone rotate around the position of the gyroscope, it actually measures the angular velocity within 1% of error. When the smartphone rotates around an off-centered axis, the measured values depend on the distance $r$ to the axis of rotation. This discrepancy originates from hardware limitations. (a) 3-axis acceleration signals recorded for a rotating phone. (b) 3-axis gyroscope signals recorder in the same rotating configuration. (c) Total gyroscopic signal $g=\sqrt{g_{x}^{2}+g_{y}^{2}}$ in this rotating configuration. (d) Measured angular velocity along $y$ from $g_{x}$ signal, as a function of the position of the smartphone with respect to the rotation axis, compared to the expected value. The imposed angular speed $\Omega$ is color-coded. e) Measurement along $g_{y}$, corresponding to a rotation around $x$ that does not physically occur. This corresponds to a systematic error, which increases linearly with $r$, and increases with rotation speed. From these measurements, we infer that the minimal error is reached when the smartphone rotates around the position of the gyroscope, we find $r_{g}=31.3\pm 1$$m$, different from the accelerometer location. f) Percentage of error $g_{y}/(\Omega r)$ as a function of the angular velocity $\Omega$, in cm-1.
The gyroscope accuracy has been tested using the rotating table introduced in section 2.1.2, the accelerometer data serving as a reference. The phone rotates along the $x$ axis at an angular frequency $\Omega=7.7$ rad/s, varying the position of the phone with respect to the rotation axis (sketch of Fig. 3(b)). Fig. 9(a) shows the average acceleration as a function of the relative sensor position $y-y_{S}$ to the rotation axis. The three colors corresponds to the components $a_{x}$, $a_{y}$ and $a_{z}$ of the acceleration. We recover that the sensor measures the acceleration of gravity along the $y$ axis for all sliding positions, the centrifugal force along $x$, and the $z$ component is identically zero. Fig. 9(b) shows the average angular velocity components $g_{x},g_{y}$ and $g_{z}$ measured simultaneously by the gyroscope, as a function of the relative smartphone position $y-y_{S}$. The main rotation is indeed found along $y$, and the value agrees quantitatively with less than 0.5% of error when the accelerometer sensor is on the rotation axis ($y=y_{s}$). However, we also measures a linear variation of $g_{x}$ as a function of $y-y_{S}$ and a small (quadratic) decrease of $g_{y}$. These measurements do not correspond to a true phone rotation as evidenced by the accelerometer signals, but rather to a limitation of the gyroscope sensor. In this specific case of a known rotation axis, the systematic error can be compensated, by noticing that $g=\sqrt{g_{x}^{2}+g_{y}^{2}}$ is conserved for all tested smartphone positions $y-y_{S}$ (Fig. 9(c)), and correspond to the imposed angular frequency $\Omega$. The same behaviour is observed for all three directions of rotation. This discrepancy originates from the limitation of the measurement technique implemented in the gyroscope sensor. In practice, the gyroscope is a vibrating structure gyroscope (VSG) which is based on the Coriolis effect applied on a vibrating mass. As a consequence, in a non Galilean frame of reference, the gyroscope measures a combination of inertial acceleration and Coriolis force, and the sensor reading deviates from the true angular frequency, both in magnitude and in direction. To further test the gyroscope sensor limitations, we perform additional measurements with increasing rotation rate, at four different phone locations $y-y_{S}$ for a rotation around the $x$ axis of the phone. Fig. 9(d) shows the main angular velocity component $g_{x}$ as a function of $y-y_{S}$, for different rotation rates (color coded). These measurements confirm that for small $\Omega$ (typically $\Omega<5$ rad/s), we recover $g_{x}\sim\Omega$. For larger rotation rate, we observe a significant discrepancy, which depends on the smartphone location. In particular for $|y-y_{S}|$ = 50 mm the error grows to 10% at $\Omega=10$ rad/s. The systematic error is best evidenced by the signals of $g_{y}$ (Fig. 9(e)), which should be identically zero. For all rotation speeds, the systematic error is linear with the distance $r$ between the sensor and the axis of rotation, and the slope increases with $\Omega$. From these measurements, we infer that the gyroscope sensor is located at the position $y_{g}=31.3\pm 1$ mm along the $y$ axis, very close to the accelerometer sensor. Eventually, we extract the slope $\alpha_{g}$ of the error $g_{y}=\alpha_{g}(r-r_{g})$, and we compute a relative error, $\Delta g=\alpha_{g}/\Omega$, which is measured in $\%/cm$, as it is an error proportional to $r$. The error is represented in Fig. 9(f) as a function of $\Omega$. For a period of 1s corresponding to $\Omega=2\pi$ rad/s, we found 4% of error per centimeter. Among the sources of error, we identify (i) mechanical coupling between the $x$ and $y$ gyroscopes, which plays a symmetrical roles. It would explain why the orientation of the rotation vector is modified but the total angular velocity $\Omega$ is conserved at moderate rotation and moderate distance to the axis of rotation. The assumption of pure Coriolis force measured by $g_{i}$ fails for increasing velocity and acceleration of the gyroscope sensors, it adds at least two other sources of error : (ii) The sensor also measure inertial forces, such as centrifugal force, which is proportional to $r$. (iii) The assumption of a purely oscillatory motion imposed to the sensor by the ship is no longer valid, the velocity of the gyro with respect to the Galilean frame of reference is no longer negligible. For a rotation at constant speed, all these sources of error are proportional to the distance $r$ to the axis, but depend differently on the angular velocity $\Omega$. For oscillatory motions and a distance $r$ to the center of rotation varying in time, these systematic sources of errors would be difficult to disentangled. We conclude that in practice, the gyroscope of the Redmi10A is reliable only to measure pure rotational motions around the gyroscope sensor, and cannot be used in general in combination with the accelerometer to decompose arbitrary superposition of translational and rotational motions. However, in some limit cases, in particular when the center of rotation is known, we expect the gyroscope to be reliable. One may refer to Fig. 9(f) to evaluate the systematic error as a function of both the angular frequency and the distance $r$ between the sensor and the axis of rotation.
Figure 10: Test of the gyroscope accuracy on a Fairphone 4. Averaged angular velocity $\Omega_{g}$ measured with the gyroscope as a function of the angular velocity $\Omega_{m}$ measured with the magnetometer using $\Omega_{m}=2\pi f_{0}$, where $f_{0}$ is the frequency of oscillation of the horizontal components of the magnetometer. The Red dashed line corresponds to a line of slope 1, showing a quantitative agreement between the two measurement techniques.
For comparison, we test another smartphone, the Fairphone 4 (FP4), which also embeds an accelerometer, a gyroscope, and a magnetometer. We conduct the same series of tests on the FP4. The smartphone is mounted on a rotating table, as sketched on figure 1b), rotating at a constant angular velocity $\Omega$. We first measure the angular velocity $\Omega_{m}$ with the magnetometer using the frequency of oscillation $f_{0}$ of any component of the magnetometer sensor. The frequency $f_{0}$ is extracted from the Fourier spectrum of the component $x$ of the magnetometer, and we have $\Omega_{m}=2\pi f_{0}$. The angular velocity is also measured from the gyroscope sensor. We measure the component of the gyroscope along the direction of the rotation axis, and we extract its temporal average $\Omega_{g}$, for different rotation speed. Figure 10 shows the angular velocity $\Omega_{g}$ measured with the gyroscrope sensor, as a function of the angular velocity $\Omega_{m}$ measured with the magnetometer. The red dashed line corresponds to a line of slope 1, and it fits perfectly the data points. Therefore, the gyroscope sensor of the FP4 measures the angular velocity without any significant error up to the maximum tested rotation speed ($\Omega\sim 10$ rad/s), independent of the distance between the sensor and the axis of rotation. We conclude that the limitation of the sensors and their accuracy can vary drastically from one smartphone to another, and a benchmark of the sensors, following for instance the protocols presented in this article, is necessary.
Appendix B Mechanical validation of the software synchronization of the smartphone clock
B.1 Mechanical wave propagation
Figure 11: Conceptual sketch for (a) time synchronization using the propagation of Lamb waves in a stiff plate. (b) Zoom view on the vertical acceleration signal after one hit at the center of the table. The response arrives on all the phone within one sampling period. c) Display of the full recording (32 hits), showing that all phones receives each vertical acceleration within two time clock difference $2t_{s}$, and smaller than one on average. d) Averaged time difference $\Delta t$ of arrival between the first phone and each of the 21 other phones. The time difference $\Delta t$ remains smaller than the sampling period $t_{s}=2.5$ms for all smartphones. The errorbar corresponds to the standard deviation of the time difference distribution.
We have conducted experiments to mechanically test the smartphone synchronization. The acceleration signals have a sampling frequency of $400$ Hz, which implies a two-step approach to verify the validity of our value. We first placed $22$ smartphones (labeled from #10\#10 to #31\#31) on a metallic plate along a circular pattern (see fig. 11a) and excited impulsional lamb waves from the center after synchronizing them through our custom UDP-based protocol. Fig 11b presents such a typical impulsion as it is recorded by the 22 synchronized smartphones, showing that all smartphones record the response within one sampling period. We further confirm this by repeating a sequence of 32 hits on the plate (Fig 11c). We can compute the averaged time difference $\Delta t$ of arrival between the first phone and each of the 21 other phones. The time difference $\Delta t$ remains smaller than the sampling period $t_{s}=2.5$ms for all smartphones. The errorbar corresponds to the standard deviation of the time difference distribution. This synchronization method, while reliable, cannot provide an estimate significantly better than the sampling period $t_{s}$.
B.2 Phase measurement of two oscillating phone
[…]
Appendix C Operational constraints of the smartphone fleet
This appendix summarizes practical engineering considerations encountered in our specific experimental implementation of the smartphone fleet, including power consumption, data storage, network communication, and data recovery.
Battery consumption Several tests were performed to evaluate battery consumption during long-duration recordings, as a function of screen state (on/off), illumination, external temperature, and smartphone protection. Both ambient temperature and screen illumination were found to significantly affect battery lifetime.
Screen state has a strong impact on battery lifetime. The most energy-efficient configuration corresponds to a screen-off setting, with short periods of activity (5min every hour), for which a periodic recording duration of 48 hours was achieved. Under identical temperature conditions, keeping the screen on reduces the maximum recording duration to 20 hours. Note that these values will strongly depend on the phone model, due to battery capability, and battery optimization softwares.
Ambient temperature also plays a critical role. When the external temperature is decreased from to -20°C, the battery lifetime of non insulated smartphone drops to approximately 8 hours. In the current configuration, the operational temperature range of a non insulated smartphone extends from 20-20 °C to $25$ °C. Finally, the use of a custom plastic protective case significantly improves low-temperature performance as their provide thermal shielding. With this enclosure, smartphones remain operational at temperatures as low as 20-20 °C, with a battery lifetime larger than 14 hours.
Data-volume management The Redmi 10A smartphone provides a total usable internal storage capacity of 64 GB. For continuous recordings simultaneously using the accelerometer (400 Hz), gyroscope (400 Hz) and magnetometer (50 Hz) at their maximum frame rate, we observe a data flux of 100 MB/hour. The smartphones thus run out of battery before their internal memory is full. Data volume does not represent a limiting factor in practice.
Packet-loss rates Packet loss was primarily observed during network communication phases, such as remote start/stop commands and data retrieval, and did not affect data acquisition itself, which is performed locally on each device. In stable network conditions, command execution and data recovery through Wi-Fi were successful for more than 95 % of the smartphones. Packet losses were mainly associated with transient Wi-Fi instabilities, particularly in outdoor environments or during simultaneous data transfers from multiple devices. These connection ruptures were solved by manual recovery when needed, and did not result in data corruption or loss of recorded signals. Running Gobannos locally on the phone ensures that all the data are recorded on the phone memory. In this configuration, we did not experience any loss of data.